OpenStudy (lvthvn):

a car moving with a velocity of 20 meters/second has 1.8 x 10^5 joules of kinetic energy. what is the mass of the car? 4.6 x 10^2 kilograms 9.0 x 10^2 kilograms 1.0 x 10^3 kilograms 1.2 x 10^3 kilograms

OpenStudy (blazeryder):

\[K.E = 0.5 \times m \times v ^{2}\] rewrite to solve for m\[m={\frac{ 2 \times K.E. }{ v ^{2} }}\]Plug in variable and solve: \[m=\frac{ 2 \times 180000 }{ 20^{2} }\]\[m=\frac{ 360000 }{ 400 }\]\[m=900\] or \[9.0 \times 10^{2}kg\] I hope this helps! =)